# Inferring Ingrained Remote Information in AC Power Flows Using Neuromorphic Modality Regime

### v1

@@ Abstract
In this paper, we infer ingrained remote information in AC power flows using spiking neural network (SNN) as edge processors for efficient coordination of power electronic converters. This work unifies power and information as a means of data normalization using a multi-modal regime in the form of spikes using energy-efficient neuromorphic processing and semantics theory. Firstly, we organize the synchronous real-valued measurements at each edge and translate them into asynchronous spike-based events to collect sparse data for training of SNN at each edge. Instead of relying on error-dependent supervised data-driven learning theory, we exploit the latency-driven unsupervised Hebbian learning rule to obtain modulation pulses for switching of power electronic converters that can now communicate among each other. Not only does this philosophy block exogenous path arrival for cyber attackers by dismissing the cyber layer, it also entails converter adaptation to system reconfiguration and parameter mismatch issues. We conclude this work by validating its energy-efficient and effective online learning performance under various scenarios in modified IEEE 14-bus system and under experimental conditions.
@@ I Introduction
Digitalization of power electronic dominated grids has brought new data-driven directions in health monitoring and coordination of ever-scaling distributed energy resources (DERs). As we move towards decentralized generation, the control and reliability of DERs demand orchestration from communication networks as an overlaying control and coordination layer to cope with the ever-increasing power demand. However, these digitization efforts not only have increased cyber attack vulnerabilities but have added more unreliability metrics in the form of latency, data dropouts, etc [1], [2]. This mandates an intrinsic communication principle that can effectively unify power & information to improve power electronic system cognition and resiliency. Since the layered hierarchy of communication networks overlaying power grids introduces complexity in modeling, Power Talk and Talkative Power technologies were developed for converters to communicate via tie-lines such that the information is tactfully modulated over the power signals [3]. The communication rate is improved by modifying the switching characteristics. Several innovations involving the co-transfer of power & data via transmission lines are inspired by this technology can be found in [4]. However, the scalability of these co-transfer technologies is still an open question.
@@
To address the scalability problem, we take inspiration from biologically-plausible neuromorphic architecture as an extension of our previous work [5] for DC systems, and use spiking neural network (SNN) at each edge for multi-modal information exchange between DERs via tie-lines in an AC system and achieve coordinated control. It is important to notify that there is a significant protocol disparity on information collection since the existing co-transfer technologies [3], [4] imply on communicating between two ends by modulation the information on power signal. On the other hand, the proposed approach facilitates inference of semantic information at both ends using the power flows between them. The multi-modal inference process is steered by some of the key semantic metrics proposed in [6], [7].
@@
Going beyond the conventional cyber-physical hierarchical control for power electronic dominated grids in Fig. 1(a), we convert dynamic grid disturbances into binary and sparse asynchronous events [8] using the local measurements, that are then introduced as spikes based data for training and adapting the SNNs deployed at each DER. Locally generated spike based data customizes SNN in an online fashion using unsupervised Hebbian learning rule that reinforces adaptation of each DER for various scenarios, potentially enhancing power system security and reliable communication. Furthermore, the adaptation features allow it to be a promising energy-efficient grid-edge intelligence tool that can be integrated in a plug-and-play manner. In summary, the proposed multi-modal co-transfer technology offers the following features:
@@
    Dynamic remote estimation with energy-efficient neuromorphic edge processors deployed at each DER. This is performed by inferring ingrained remote information from the AC power flow itself at each edge. To train the SNN, the real-valued measurements are converted into asynchronous spikes for data normalization using a modality of power and information embedded into the timing notion of a spike occurrence. We exploit unsupervised Hebbian learning based rules to update the weights of SNN at each edge. As a result, SNNs can be incrementally deployed at each node as a decentralized secondary controller that significantly minimizes numerous hierarchical steps. It completely dismisses the cyber layer for system operation, such that it bypasses the cyber unreliability concerns and attack vulnerabilities altogether.
@@
    Since it replaces the conventional cyber-physical hierarchical stages for control, it also reinforces adaptive behavior into each converter as an edge facility during scenarios, such as system reconfiguration and parameter mismatch issues.
  3. 3.
    The proposed approach is an energy-efficient digital grid-edge technology that only operates during events.
@@ II Design Preliminaries of Multi-Modal Power in AC Systems
As shown in Fig. 1(a), in comparison with the conventional cyber-physical control framework for power electronic dominated grids, ingrained remote information in AC power flow can instill a very simplified architecture for each DER, which has been shown in Fig. 1(b). However, we need grid-edge tools that can effectively infer the ingrained information and exploit it for coordinated control of power electronic dominated grids. This is where we deploy a spiking neural network at each edge so that we can instill an event-based selective response and coordination. More specifically, the SNNs at each edge will update their weights in accordance with the system dynamics. Although the dynamic change is global, the impact on DERs might be local, which then condenses it to local governing equations for generation of spikes at each DER. Since it equips a spike-to-spike based SNN control strategy for each DER, it overrules many hierarchical stages from the conventional cyber-physical control framework in Fig. 1(a). Hence, we exploit a semantic sampling policy to selectively capture the most significant data corresponding to dynamic disturbances only for each DER. These analog voltage and current data are then transformed into spikes through the encoding layer, as shown in Fig. 1(c).
@@
The learning activation in SNN is performed by the leaky integrate-and-fire (LIF) neuron model that learns by performing spatio-temporal integration of the synpatic inputs, as shown in Fig. 2. Approximating the LIF neuron dynamics using the Euler method, we get:
$\tau_{m}\frac{\mathrm{d}V_{mem}}{\mathrm{d}t}=-(V_{mem}-V_{th})+\frac{I_{s}}{g_{L}}$ (1)
@@
where, $V_{mem}$ is the membrane potential, $g_{L}$ is the leaky conductance of each membrane, $V_{th}$ is the synaptic threshold, $I_{s}$ is the synaptic current and $\tau_{m}$ is the membrane time constant. Similar to a switched RC network with the same voltage potentials, each neuron can be programmed with a certain stored charge $q$ that has a corresponding notion of leakage, which only gets activated beyond a given voltage threshold $V_{th}$. The equivalence with (1) for a switched RC network can be given by:
$\tau_{RC}\frac{\mathrm{d}q}{\mathrm{d}t}=-\frac{V_{mem}-V_{th}}{C}+{q}$ (2)
@@
where, $\tau_{RC}=RC$. As it can be seen in Fig. 2, an output spike is generated only when the threshold voltage $V_{th}$ is reached. Hence, the SNNs deployed at each DER to realize becomes a sparse and energy-efficient alternative, owing to the fact the LIF neurons are active only when: (a) input events are present, (b) culmination of the input events goes beyond the activation potential $V_{th}$.
@@
Figure 2: Leaky integrate and fire (LIF) model of SNN: (a) Input and output of the SNN at each DER, where only some LIF neurons are active subject to the frequecy of spikes, (b) the neuron model that synthesizes input spikes into selective output spikes based on a voltage threshold $V_{th}$.
@@
This approach enables the DER to dynamically adapt its operation in response to various disturbances. It is worth notifying that during steady-state conditions, both semantic sampling in the input and the SNN remain idle, consuming minimal energy.
II-A Semantic Sampling
@@
As data collection is carried out in a structured way, we formalize only the most significant data to be triggered as binary events, that correspond to any physical disturbance in this paper. For instance, the event $\Omega_{vd}$, associated with the $d$-axis component $v_{d}$ of sampled voltage (see Algorithm 1), can be detected using a moving window approach using:
$\Omega_{vd}:\frac{1}{N_{w}}\sum{(\mathbf{v_{dw}}-\mathbf{\bar{v}_{dw}})^{2}}>\sigma^{V_{d}}_{th}$ (3)
@@
Remark II: It is worthy notifying that a moving window approach for event detection and spike generation has been adopted in this work to achieve a good balance between implementation efficiency, versatility and biophysical accuracy of the neurons [9]. The data movement rather assists the solver-based digital implementation tasks.
Algorithm 1 Event-driven SNN activation for Inverter $k$
@@
$m^{\mathrm{th}}$samples of system state ($dq$voltage & current) $\mathbf{x}=[\mathbf{v_{dq}}\;\mathbf{i_{dq}}]^{\mathbf{T}}$, length of the moving window $N_{w}$, and thresholds of voltage & current variance to trigger an event $\sigma^{V_{dq}}_{th}$, $\sigma^{I_{dq}}_{th}$.
$\mathbf{x_{w}}\leftarrow\{\mathbf{x}[m-N_{w}],...,\mathbf{x}[m-1]\}$, $\mathbf{\bar{x}_{w}}=\frac{1}{N_{w}}\sum{\mathbf{x_{w}}}$
@@
5:Transfer $\mathbf{x}_{k}[m]$ to SNN
6:Update $\mathbf{x_{w}}$ and $\mathbf{\bar{x}_{w}}$ by the $m$-th sample:
7:$\mathbf{x_{w}}\leftarrow\{\mathbf{x}[m-N_{w}+1],...,\mathbf{x}[m]\}$, $\mathbf{\bar{x}_{w}}=\frac{1}{N_{w}}\sum{\mathbf{x_{w}}}$
8:$m\leftarrow m+1$
@@
When a dynamic change occurs in $v_{d}$, $\Omega_{vd}$ is triggered to a transition from 0 to 1, and returns to 0 when $v_{d}$ reaches a new steady state. A similar approach is applied to detect dynamics in $v_{q}$, $i_{d}$, and $i_{q}$, denoted as $\Omega_{vq}$, $\Omega_{id}$, and $\Omega_{iq}$, respectively. Semantic sampling [6] and sparse data collection is subsequently triggered by $\Omega$, which can be denoted by:
$\Omega=\Omega_{vd}\quad\texttt{OR}\quad\Omega_{vq}\quad\texttt{OR}\quad\Omega_{id}\quad\texttt{OR}\quad\Omega_{iq}$ (4)
Using (4), only when $\Omega$ is 1, the corresponding dynamics in the measurements $\{v_{k},i_{k}\ \forall k\in N_{\mathrm{DER}}\}$ of inverter $k$ will be sampled for inputs to the inferential learning of SNN, as described in Algorithm 1.
@@ II-B1 Encoding of Input Samplings
As depicted in Fig. 1(c), we employ rate encoding method to convert the real-valued samples $v_{k}$ and $i_{k}$ into spikes. In rate encoding, a higher input value results in a greater number of spikes within a specific time interval. For instance, considering $i_{a}$ as an example in Fig. 1(c), spikes are generated only when $i_{a}[m]$ exceeds the carrier wave.
@@ II-B2 Training of SNN
In Fig. 1(c), the input to SNN comprises spikes originating from the encoding layer, which encapsulates crucial dynamic information. The final objective behind the training process is to shape these spikes into a PWM signal. To achieve this, we utilize modulation signals $\bm{m_{abc}}$ in Fig. 1(a) as the regression target to initiate SNN, obtained from the control of virtual synchronous generator (VSG). With reactive power sharing and average voltage regulation being the coordination objectives of the secondary controller, fault ride-through (FRT) capability [11] is seamlessly integrated into the SNN. Since $\bm{m_{abc}}$ is an analog signal, we employ the membrane potential $\bm{u_{abc}}$ of three neurons in the output layer to align with it. Training the SNN involves the well-known back-propagation algorithm in conjunction with the mean squared error (MSE) loss function:
$\texttt{Loss}=\frac{1}{N_{w}}\sum_{i=1}^{N}(\bm{m_{abc}}-\bm{u_{abc}})^{2}$ (5)
@@ II-B3 Decoding of SNN
To convert the membrane potential $\bm{u_{abc}}$ into spikes for PWM generation, we employ a dedicated threshold signal, designed to match the shape of the carrier wave as shown in Fig. 1 using:
$\begin{array}[]{rr}{\texttt{Threshold}=\cfrac{2}{T_{s}}\left[t-\cfrac{2N-1}{2}T_{s}\right]}&{(N-1)T_{s}\leq t<NT_{s}}\\ \end{array}$ (6)
where, $T_{s}$ represents the carrier signals’ period. We adapt the membrane potential model from [10] by eliminating the influence of output spikes. When the membrane potential surpasses the Threshold in (6), spikes are generated. This neuron model aligns with the modulation process of the sinusoidal pulse width modulation (SPWM). Thus, if the membrane potential accurately matches $\bm{m_{abc}}$, the output spikes will also map the PWM signals accordingly.

### v2

@@ Abstract
In this paper, we infer remote measurements such as remote voltages and currents online with change in AC power flows using spiking neural network (SNN) as grid-edge technology for efficient coordination of power electronic converters. This work unifies power and information as a means of data normalization using a multi-modal regime in the form of spikes using energy-efficient neuromorphic learning and event-driven asynchronous data collection. Firstly, we organize the synchronous real-valued measurements at each edge and translate them into asynchronous spike-based events to collect sparse data for training of SNN at each edge. Instead of relying on error-dependent supervised data-driven learning theory, we exploit the latency-driven unsupervised Hebbian learning rule to obtain modulation pulses for switching of power electronic converters that can now comprehend grid disturbances locally and adapt their operation without requiring explicit infrastructure for global coordination. Not only does this philosophy block exogenous path arrival for cyber attackers by dismissing the cyber layer, it also entails converter adaptation to system reconfiguration and parameter mismatch issues. We conclude this work by validating its energy-efficient and effective online learning performance under various scenarios in different system sizes, including modified IEEE 14-bus system and under experimental conditions.
@@ I Introduction
Digitalization of power electronic dominated grids has brought new data-driven directions in health monitoring and coordination of ever-scaling distributed energy resources (DERs). As we move towards decentralized generation, the control and reliability of DERs demand orchestration from communication networks as an overlaying control and coordination layer to cope with the ever-increasing power demand. However, these digitization efforts not only have increased cyber attack vulnerabilities but have added more unreliability performance in the form of latency, data dropouts, etc [1], [2].
To this end, many cyber security detection and mitigation technologies for power electronic dominated grids have surfaced that can be augmented easily into the cyber [3] as well as the physical layer [4]. However, the pace of evolving cyber-physical vulnerabilities can outgrow these research developments quite easily. Apart from security concerns, information exchange enabled by the cyber layer also faces reliability concerns with communication network traffic leading to latency, data dropouts, link failure, etc. As a result, the stability and reliability of power electronic dominated grids is affected [1], [5]. Although decentralized coordination using droop control strategy [6] addresses the said concerns in a monolithic manner, its error-immanent operation affects the system performance and efficiency. This mandates an intrinsic communication principle that can effectively unify power & information to improve power electronic system cognition and resiliency.
Since the layered hierarchy of communication networks overlaying power grids introduces complexity in modeling, Power Talk and Talkative Power technologies were developed for converters to communicate via tie-lines such that the information is tactfully modulated over the power signals [7]. The communication rate is improved by modifying the switching characteristics. Several innovations involving the co-transfer of power & data via transmission lines [^1] They are commonly referred to as tie-lines. Both terms are interchangeably used in this paper as per the context. are inspired by this technology can be found in [8]. However, the scalability of these co-transfer technologies is still an open question.
@@
To address the scalability problem, we take inspiration from biologically-plausible neuromorphic architecture as an extension of our previous work [9], [10] for DC systems, and use spiking neural network (SNN) at each edge for multi-modal information exchange between DERs via tie-lines in an AC system and achieve coordinated control. The multi-modal inference scheme is steered by some of the key event-driven semantic metrics proposed in [11], [12].
@@
Going beyond the conventional cyber-physical hierarchical control for power electronic dominated grids in Fig. 1(a), we convert dynamic grid disturbances into binary and sparse asynchronous events [13] using local measurements, that are then introduced as spikes based data for online training and adapting the SNNs deployed at each DER. Locally generated spikes customize SNN at each edge in an online fashion using unsupervised Hebbian learning rule that comprehends the power grid network disturbances locally, and adapts DERs’ response. Furthermore, the adaptation features allow it to be a promising energy-efficient grid-edge intelligence tool that can be integrated in a plug-and-play manner. In summary, the proposed technology offers the following features:
@@
    Dynamic remote estimation with energy-efficient neuromorphic event-driven grid-edge processor deployed at each DER. This is performed by inferring ingrained remote information from the AC power flow itself at each edge. To train the SNN, the real-valued measurements are converted into asynchronous spikes for data normalization using a modality of power and information embedded into the timing notion of a spike occurrence. We exploit unsupervised Hebbian learning based rules to update the weights of SNN at each edge. As a result, SNNs can be incrementally deployed at each node as a decentralized secondary controller that significantly reduces hierarchical control stages (see Fig. 1(b)). By completely dismissing the cyber layer for system operation, it bypasses the cyber unreliability concerns and attack vulnerabilities altogether.
@@
    Since it replaces the conventional cyber-physical hierarchical stages for control, it also reinforces adaptive behavior into each DER to reconfigure its operation during scenarios, such as system reconfiguration and parameter mismatch issues.
@@ II Design Preliminaries of Multi-Modal Power in AC Systems
As shown in Fig. 1(a), in comparison with the conventional cyber-physical control framework for power electronic dominated grids, ingrained remote information in AC power flow can instill a very simplified architecture for each DER, which has been shown in Fig. 1(b). However, we need grid-edge tools that can effectively comprehend remote information, such as remote voltages, currents, power and exploit it for coordinated control of power electronic dominated grids. This is where we deploy a spiking neural network (SNN) at each edge so that we can instill an event-based selective data collection. More specifically, the SNNs at each edge will update their weights in accordance with dynamic conditions in the system. Although the dynamic change is global, the impact on DERs might be local, which then condenses it to local governing equations for generation of spikes at each DER. Since it equips a spike-to-spike based SNN control strategy for each DER, it overrules many hierarchical stages from the conventional cyber-physical control framework in Fig. 1(a). As shown in Fig. 1(c), we exploit an event-driven sampling policy to selectively capture the most significant data corresponding to dynamic disturbances only. These analog voltage and current data are then transformed into spikes through the encoding layer, as shown in Fig. 1(c). For example, it can be seen in Fig. 1(c) that spikes are only generated as per the event-driven sampling criteria only when a dynamic change is encountered in the dq signals.
@@
The learning activation in SNN is performed by the biologically plausible leaky integrate-and-fire (LIF) neuron model that learns by performing spatio-temporal integration of the synaptic inputs, as shown in Fig. 2. Approximating the LIF neuron dynamics using the Euler method, we get:
$\mathrm{\tau_{m}\frac{\mathrm{d}V_{mem}}{\mathrm{d}t}=-(V_{mem}-V_{th})+\frac{I_{s}}{g_{L}}}$ (1)
@@
where, $\mathrm{V_{mem}}$ is the membrane potential, $\mathrm{g_{L}}$ is the leaky conductance of each membrane, $\mathrm{V_{th}}$ is the synaptic threshold, $\mathrm{I_{s}}$ is the synaptic current and $\mathrm{\tau_{m}}$ is the membrane time constant. Similar to a switched RC network with the same voltage potential, each neuron can be programmed with a certain stored charge $\mathrm{q}$ that has a corresponding notion of leakage, which only gets activated beyond a given voltage threshold $\mathrm{V_{th}}$. The equivalence with (1) for a switched RC network can be given by:
$\mathrm{\tau_{RC}\frac{\mathrm{d}q}{\mathrm{d}t}=-\frac{V_{mem}-V_{th}}{C}+{q}}$ (2)
@@
where, $\mathrm{\tau_{RC}=RC}$. As it can be seen in Fig. 2, an output spike is generated only when the threshold voltage $\mathrm{V_{th}}$ is reached. Hence, the SNN becomes a sparse and energy-efficient alternative, owing to the fact the LIF neurons are active only when: (a) input events are present, (b) culmination of the input events goes beyond the activation potential $\mathrm{V_{th}}$.
@@
Figure 2: Leaky integrate and fire (LIF) model of SNN: (a) Input and output of the SNN at each DER, where only some LIF neurons are active subject to the frequency of spikes, (b) the neuron model that synthesizes input spikes into selective output spikes based on a voltage threshold $\mathrm{V_{th}}$.
@@
This approach enables the DER to dynamically adapt its operation in response to various disturbances. It is worth notifying that during steady-state conditions, both event-driven sampling in the input and the SNN remain idle, consuming minimal energy.
II-A Event-Driven Sampling
@@
To carry out a structured and sparse mechanism for data collection, we formalize only the most significant data to be triggered as binary events, that correspond to any physical disturbance in this paper. For instance, the event $\mathrm{\Omega_{vd}}$, associated with the $\mathrm{d}$-axis component $\mathrm{v_{d}}$ of sampled voltage (see Algorithm 1), can be detected using a moving window approach using:
$\mathrm{\Omega_{vd}:\frac{1}{N_{w}}\sum{(\mathbf{v_{dw}}-\mathbf{\bar{v}_{dw}})^{2}}>\sigma^{V_{d}}_{th}}$ (3)
where, $\mathrm{\mathrm{N_{w}}}$ is the length of the moving window.
@@
Remark: It is worthy notifying that a moving window approach for event detection and spike generation has been adopted in this work to achieve a good balance between implementation efficiency, versatility and biophysical accuracy of the neurons [14]. The data movement rather assists the solver-based digital implementation tasks.
Algorithm 1 Event-driven SNN activation for Inverter $\mathrm{k}$
@@
$\mathrm{m^{\mathrm{th}}}$samples of system state (dqvoltage & current) $\mathrm{\mathbf{x}=[\mathbf{v_{dq}}\;\mathbf{i_{dq}}]^{\mathbf{T}}}$, length of the moving window $\mathrm{N_{w}}$, and thresholds of voltage & current variance to trigger an event $\mathrm{\sigma^{V_{dq}}_{th}}$, $\mathrm{\sigma^{I_{dq}}_{th}}$.
$\mathrm{\mathbf{x_{w}}\leftarrow\{\mathbf{x}[m-N_{w}],...,\mathbf{x}[m-1]\}}$, $\mathrm{\mathbf{\bar{x}_{w}}=\frac{1}{N_{w}}\sum{\mathbf{x_{w}}}}$
@@
5:Transfer $\mathrm{\mathbf{x}_{k}[m]}$ to SNN
6:Update $\mathrm{\mathbf{x_{w}}}$ and $\mathrm{\mathbf{\bar{x}_{w}}}$ by the $\mathrm{m}$-th sample:
7:$\mathrm{\mathbf{x_{w}}\leftarrow\{\mathbf{x}[m-N_{w}+1],...,\mathbf{x}[m]\}}$
8:$\mathrm{\mathbf{\bar{x}_{w}}=\frac{1}{N_{w}}\sum{\mathbf{x_{w}}}}$
9:$\mathrm{m\leftarrow m+1}$
@@
When a dynamic change occurs in $\mathrm{v_{d}}$, $\mathrm{\Omega_{vd}}$ is triggered to a transition from 0 to 1, and returns to 0 when $\mathrm{v_{d}}$ reaches a new steady state. A similar approach is applied to detect dynamics in $\mathrm{v_{q}}$, $\mathrm{i_{d}}$, and $\mathrm{i_{q}}$, denoted as $\mathrm{\Omega_{vq}}$, $\mathrm{\Omega_{id}}$, and $\mathrm{\Omega_{iq}}$, respectively. Event-driven sampling [11] and sparse data collection is subsequently triggered by $\mathrm{\Omega}$, which can be given by:
$\mathrm{\Omega=\Omega_{vd}\quad\texttt{OR}\quad\Omega_{vq}\quad\texttt{OR}\quad\Omega_{id}\quad\texttt{OR}\quad\Omega_{iq}}$ (4)
Using (4), only when $\mathrm{\Omega}$ is 1, the corresponding dynamics in the measurements $\mathrm{\{v_{k},i_{k}\ \forall\ k\in N_{\mathrm{DER}}\}}$ of inverter $\mathrm{k}$ will be sampled for inputs to the inferential learning of SNN, as described in Algorithm 1.
@@ II-B1 Encoding of Input Samplings
As depicted in Fig. 1(c), we employ rate encoding method to convert the real-valued samples $\mathrm{v_{k}}$ and $\mathrm{i_{k}}$ into spikes. In rate encoding, a higher input value results in a greater number of spikes within a specific time interval. For instance, considering $\mathrm{i_{a}}$ as an example in Fig. 1(c), spikes are generated only when $\mathrm{i_{a}[m]}$ exceeds the carrier wave.
@@ II-B2 Training of SNN
In Fig. 1(c), the input to SNN comprises spikes originating from the encoding layer, which encapsulates crucial dynamic information. The final objective behind the training process is to shape these spikes into a PWM signal. To achieve this, we utilize modulation signals $\mathrm{\mathbf{m_{abc}}}$ in Fig. 1(a) as the regression target to initiate SNN, obtained from the control of virtual synchronous generator (VSG). Since $\mathrm{\mathbf{m_{abc}}}$ is an analog signal, we employ the membrane potential $\mathrm{\mathbf{u_{abc}}}$ (see Fig. 1(c)) of three neurons in the output layer to align with it. Training the SNN involves the well-known back-propagation algorithm in conjunction with the mean squared error (MSE) loss function:
$\mathrm{\texttt{Loss}=\frac{1}{N_{w}}\sum_{i=1}^{N}(\mathbf{m_{abc}}-\mathbf{u_{abc}})^{2}}$ (5)
@@ II-B3 Decoding of SNN
To convert the membrane potential $\mathrm{\mathbf{u_{abc}}}$ into spikes for PWM generation, we employ a dedicated threshold signal, designed to match the shape of the carrier wave as shown in Fig. 1 using:
$\mathrm{\texttt{Threshold}=\cfrac{2}{T_{s}}\left[t-\cfrac{2N-1}{2}T_{s}\right]},\;\mathrm{(N-1)T_{s}\leq t<NT_{s}}$ (6)
where, $\mathrm{T_{s}}$ represents the carrier signals’ period. We adapt the membrane potential model from [15] by eliminating the influence of output spikes. When the membrane potential surpasses the Threshold in (6), spikes are generated. This neuron model aligns with the modulation process of the sinusoidal pulse width modulation (SPWM). Thus, if the membrane potential accurately matches $\mathrm{\mathbf{m_{abc}}}$, the output spikes will also map the PWM signals accordingly.